elimination method
1.3x-0.2y=12
0.4+17y=89
show work please

Answer by OHtutor(16)

multiply 1st equation by 10, and 2nd one as well.
13x-2y=120
4+170y=890
y=(890-4)/170=786/170=393/85
13x=120+2y=120+2.393/85=(120.85+786)85
I guess you can continue now.
Where did u get those number from? Very weird....